# Laplace measure transitions and ghosts for meromorphic functions

### 05 Apr 2026 v1

@@ Abstract
We study the measure transition problem for bilateral Laplace transforms of meromorphic functions on vertical strips. Given a meromorphic function $F$ admitting Laplace representations on two adjacent strips separated by a vertical line, we investigate how the corresponding determining measures are related. Our first result shows that in the absence of poles on the separatrix the determining measures coincide. We next derive explicit transition formulas for the case of finitely many poles and obtain sufficient conditions under which these formulas remain valid for infinitely many poles. Applications are given to the analytic continuation of the zeta function, periodic and almost periodic functions, and quotients of Gamma functions related to the confluent hypergeometric function. Finally, using generalized Cauchy integrals, we construct an entire function admitting distinct Laplace representations on the right and left half-planes, thereby producing a ghost transition. This provides a counterexample to uniqueness of solutions of the Cauchy problem for the heat equation.
@@ 1 Introduction
This function is holomorphic in its maximal strip of convergence
$\Omega_{a,b}=\{s\in\mathbb{C}:a<\Re(s)<b\},$
where $\infty\leq a<b\leq+\infty.$
@@
Determine the relationship between the determining measures $\mu_{r}$ and $\mu_{l}.$
Several approaches to the measure transition problem have appeared in the literature. The first mention of the problem appears to date back to Van der Pol [31], who used classical Laplace inversion techniques to derive a transition formula for the difference $\mu_{r}-\mu_{l}$ in the case of a finite number of poles. His method relied on the uniform vanishing of the quotient $F(s)/s$ as $|\Im(s)|\to+\infty$ inside the strip $\Omega_{a,b}$. The significance of the problem was demonstrated through applications of the transition formula to the analytic continuation of several meromorphic functions of number-theoretic interest, including the Riemann zeta function. The subject was later revisited independently by Buescu et al. [2, 3, 4, 5, 6] and Harper [14]. In the first case the context is that of polar (co-positive or co-definite, see definition 2.5 below) meromorphic functions. In that setting, the determining measures are decomposed into principal and holomorphic parts, and a transition formula for $d\mu_{r}-d\mu_{l}$ was obtained under the assumption that the holomorphic components extend across the separatrix. A different point of view was considered by Harper, who established a connection between the existence of two Laplace representations for a holomorphic function $F$ on a vertical strip and solutions of the Cauchy problem for the heat equation on the infinite rod. By imposing suitable bounds on the $L^{2}$-norm of $F$ along vertical lines, Harper applied uniqueness results for the heat equation to obtain sufficient conditions under which $\mu_{r}=\mu_{l}$.
@@ 2 Notation and definitions
The classical bilateral Laplace integral
$\int_{-\infty}^{+\infty}e^{-st}f(t)\,dt$
of a locally integrable real-valued function of a real variable $f(t)$ can be seen as a special case of (1) by setting $\mu(t)=\int_{0}^{t}f(u)\,du.$ Note that this definition implies that $\mu(t)$ is normalized, and we sometimes also refer to $f(t)$ and $(f_{l},f_{r})$ as normalized determining function and Laplace pair respectively. The usage will always be clear from the context.
@@
Note that co-positive definite functions correspond to classical positive definite functions on vertical lines within the strip; along each such lines Widder’s characterization (2) reduces to Bochner’s theorem (see [26] for details).
@@
Let $F(s)=\int_{-\infty}^{+\infty}e^{-st}\,d\mu(t)$ have a vertical strip of convergence $\Omega_{a,b},{\rm with}-\infty<a<b<+\infty$. For every $\epsilon>0$ and $0<\delta<\frac{|b-a|}{2},$ there exists a positive constant $Y$ such that for all $a+\delta\leq x\leq b-\delta$
$\Bigg|\frac{F(x+iy)}{y}\Bigg|<\epsilon,$
whenever $|y|>Y.$
@@
We will need the version of the classical Phragmén-Lindelöf theorem for a vertical half-strips.
@@
Let $D$ be a vertical half-strip defined by those points $a<\Re(s)<b$ lying either above or below the line $\Im(s)=d.$ Let $F(s)$ be holomorphic in $D$ and continuous up to its boundary. Assume that $|F(s)|\leq M$ on boundary lines. If there exist positive constants $A,$ $B$ and $0<K<\frac{\pi}{|b-a|}$ such that for all $s$ inside $D$ we have
$|F(x+iy)|\leq Ae^{Be^{K|y|}},$ (4)
@@ 3 The transition formula
From this point onwards we suppose the function $F(s)$ is represented by a Laplace pair $(\mu_{l},\mu_{r})$ on a vertical strip $\Omega_{a,b}$ with $a<0<b$, where the imaginary axis $\Re(s)=0$ is the separatrix. This choice is notationally convenient and clearly implies no loss of generality, since a Laplace pair with separatrix $\Re s=c$ reduces to this setting after a translation of the determining measures by the factor $e^{-ct}$.
@@
Let $F(s)$ be a holomorphic Laplace pair $(\mu_{l},\mu_{r})$ satisfying the Phragmén-Lindelöf condition 2.10 and having no singularities on the separatrix. Let $0<\delta<\min\{|a|,|b|,1\}.$ Then there exist positive constants $R_{0}$ and $M$ such that
@@
By Lemma 2.8 it follows that $F(s)/s$ is bounded by some constant $M_{\delta}>0$ on the vertical lines $|\Re(s)|=\delta.$
Since $F(s)/s$ satisfies the Phragmén–Lindelöf condition, we may choose $R_{0}>0$ such that the bound (5) holds for $F(s)/s$ whenever $|y|\geq R_{0}.$
@@
We consider the contour integral
$\frac{1}{2\pi i}\int_{\Gamma_{R,\delta}}e^{st}\frac{F(s)}{s}\,ds.$ (7)
To prove the result it suffices to show that for a suitable choice of $\delta,$ the contributions of the upper and lower horizontal edges of $\Gamma_{R,\delta}$ become arbitrarily small for all sufficiently large $R.$ Indeed, under this assumption, the inversion formula (3) together with the residue theorem yields
$\lim_{R\to+\infty}\frac{1}{2\pi i}\int_{\Gamma_{R,\delta}}e^{st}\frac{F(s)}{s}\,ds=\mu_{r}(t)-\mu_{l}(t)+\mu_{l}(+\infty)-\mu_{r}(-\infty)=F(0)$ (8)
@@
As stated in the Introduction, the inversion argument used by Van der Pol (see [31], Chapter VI, section 12) rested on the assumption that $F(s)/s$ vanishes uniformly to zero inside the whole strip $\Omega_{a,b}$ as $|\Im(s)|\to+\infty.$ The Phragmén-Lindelöf condition 2.10 is the precise analytic mechanism that guarantees that this occurs. As will be shown in Section 6, if condition 2.10 is relaxed, the growth of $F(s)$ on the separatrix may prevent the uniform vanishing of the horizontal edges of the contour, producing examples where $\mu_{r}\neq\mu_{l}.$
The following Corollary generalizes Van der Pol’s transition formula to the case of finite poles of any order.
@@
Let $F(s)$ be a meromorphic Laplace pair $(\mu_{l},\mu_{r})$ having finitely many poles $s=ip_{n}$, $1\leq n\leq N$, on the imaginary axis, and satisfying the Phragmén–Lindelöf condition 2.10. Then for each real $t$,
@@
Proceeding as in the proof of Theorem 3.2, the estimates (6) and (9) remain valid since none of the half-strips contain any poles by Remark 2.11. Therefore the same contour argument applies with the only modification that the poles $s=ip_{n}$ now lie inside the contour.
By the residue theorem,
@@
Subtracting this identity from (11) yields (10) as claimed. ∎
@@
Let $F(s)$ be a meromorphic Laplace pair $(f_{l},f_{r})$ satisfying the hypotheses of Corollary 3.3. For each pole $s=ip_{n}$, let $r_{n}$ denote its order and let
$P_{n}(s)=\sum_{k=1}^{r_{n}}\frac{a_{n,k}}{(s-ip_{n})^{k}}$
be the principal part of its Laurent expansion around $s=ip_{n}$. Then
$f_{r}(t)-f_{l}(t)=\sum_{n=1}^{N}e^{tp_{n}}\sum_{k=1}^{r_{n}}a_{n,k}\frac{t^{k-1}}{(k-1)!}.$ (12)
@@
$e^{st}\frac{F(s)}{s}=e^{itp_{n}}\left[\sum_{\begin{subarray}{c}l,m\geq 0\\ k\geq-r_{n}\end{subarray}}\frac{(-1)^{m}}{(ip_{n})^{m+1}}\frac{t^{l}}{l!}b_{n,k}(s-ip_{n})^{m+l+k}\right],$ (13)
@@
Finally, summing over all poles yields (12) which completes the proof ∎
@@ 4 The case of infinitely many poles
To establish the Laplace representation for $\Re(s)>0$, it suffices to show that for every $\epsilon>0$ there exists $C_{\epsilon}>0$ such that for all $t\geq 0$
$|f(t)|\leq C_{\epsilon}e^{\epsilon t}.$
For $t\geq 0$, we have
$\frac{t^{k-1}}{(k-1)!}\leq\frac{e^{\epsilon t}}{\epsilon^{k-1}},$
and therefore
@@
where $C_{\epsilon}<\infty$ by (17).
This bound justifies the interchange of summation and integration, showing that $f(t)$ is a locally integrable. For $\Re(s)>0$,
@@
Let the poles of $F(s)$ be such that its principal part $P(s)$ satisfies the hypotheses of Lemma 4.1. If the holomorphic part of $F$, defined by
H(s)=F(s)P(s)H(s)=F(s)-P(s)
satisfies the Phragmén–Lindelöf condition (5), then
@@
Since H(s)=F(s)P(s)H(s)=F(s)-P(s) is holomorphic and satisfies the Phragmén–Lindelöf condition, Theorem 3.2 implies $\mu_{r,H}(t)=\mu_{l,H}(t)$ and therefore
@@
where $f(t)$ is given by (16).
Estimate (18) ensures that the series defining $f(t)$ converges uniformly on compact sets, so integration may be carried out term by term. Using the residue computations (14) and (15) in Lemma 3.4 for the finite pole case, we obtain the transition formula (19) as desired. ∎
@@ 5 Examples and applications
$\mu_{l}(t)=\begin{cases}\frac{1}{2}-\{e^{t}\},\,t>0,\,e^{t}\notin\mathbb{Z}\\ n+1-e^{t},\,t=\log(n+1),\,n\geq 0\\ \frac{1}{2}-e^{t},\,t<0\end{cases}$ (22)
It is not difficult to verify that $\mu_{l}$ has $\Omega_{0,1}$ as its strip of convergence. Hence on that region
@@
We remark that in [5] (see section 4.5), the authors proved that $\zeta$ must lose its polarity upon crossing the separatrix $\Re(s)=1.$ Our explicit representation on $\Omega_{0,1}$ provides another proof of that fact since $\mu_{0,1}$ is not monotonic, and therefore the corresponding measure is neither positive nor negative, implying that $F$ in neither co-positive definite nor co-negative definite.
Since there are no poles on the imaginary axis, estimate (21) allows us to prove that the zeta function does not have a Laplace representation on any strip $\Omega_{a,0}$ adjacent to the left of the imaginary axis. Indeed, if that were the case, direct application of Theorem 3.2 shows that its determining measure would necessarily be equal to $\mu_{0,1}$. But this is impossible since $\Omega_{0,1}$ is the maximal strip of convergence of $\mu_{0,1}$.
@@
Observe that (24) and (26) recover the classical Laplace representations of $\frac{\zeta(s)}{s}$ on successive vertical strips of holomorphy as described in Titchmarsh, Chapter III[29].
@@
Let $f(t)$ be a bounded periodic function of period $T$. Its Laplace transform
@@
The corresponding determining function is the $\pi$-periodic function $f(t)=\operatorname{sgn}(\sin t),$ whose Fourier series is given by
$\frac{2}{i\pi}\sum_{n\in\mathbb{Z}}\frac{1}{2n+1}e^{i(2n+1)t}.$
While this series does not converge absolutely, it is not difficult to show that its Cèsaro means converge to $f$ in $L^{1}.$ Moreover, this is always true for bounded periodic functions. Therefore, when $f(t)$ is a bounded periodic determining function, the transition formula (27) can always be interpreted as converging in $L^{1}$ via Cèsaro means.
@@
Example 5.4 (A quotient of Gamma Functions).
@@
$_{1}F_{1}(a:b;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}}{(b)_{n}}\frac{z^{n}}{n!}\qquad(a)_{n}=\frac{\Gamma(a+n)}{\Gamma(a)},$
@@
Stirling’s formula for the Gamma function (see [25], 2.1.8) shows that

### 07 May 2026 v2

@@ Abstract
We study the measure transition problem for bilateral Laplace transforms of meromorphic functions on vertical strips. Given a meromorphic function $F$ admitting Laplace representations on two adjacent strips separated by a vertical line, we investigate how the corresponding determining measures are related. Our first result shows that in the absence of poles on the separatrix the determining measures coincide. We next derive explicit transition formulas for the case of finitely many poles and obtain sufficient conditions under which these formulas remain valid for infinitely many poles. Applications are given to the analytic continuation of the Riemann $\zeta$ function, periodic and almost periodic functions, and quotients of $\Gamma$ functions related to the confluent hypergeometric function. Finally, using generalized Cauchy integrals, we construct an entire function admitting distinct Laplace representations on the right and left half-planes, thereby producing a ghost transition. This provides a counterexample to uniqueness of solutions of the Cauchy problem for the heat equation.
@@ 1 Introduction
Denote by $\Omega_{\alpha,\beta}=\{s\in\mathbb{C}:\alpha<\Re(s)<\beta\}$, where $\infty\leq\alpha<\beta\leq+\infty$, the vertical strip on the complex plane bounded by $\alpha,\beta$. Classical Laplace transform properties ensure that $\mathcal{L}(\mu)(s)$ is holomorphic in its maximal strip of convergence $\Omega_{a,b}$, with $\infty\leq a<b\leq+\infty$.
@@
Establish the relationship between the determining measures $\mu_{r}$ and $\mu_{l}.$
Several approaches to the measure transition problem have appeared in the literature. The first mention of the problem appears to date back to Van der Pol [31], who used classical Laplace inversion techniques to derive a transition formula for the difference $\mu_{r}-\mu_{l}$ in the case of a finite number of poles. His method relied on the uniform vanishing of the quotient $F(s)/s$ as $|\Im(s)|\to+\infty$ inside the strip $\Omega_{a,b}$. The significance of the problem was demonstrated through applications of the transition formula to the analytic continuation of several meromorphic functions of number-theoretic interest, including the Riemann $\zeta$ function. The subject was later revisited independently by Buescu et al. [2, 3, 4, 5, 6] and Harper [14]. In the first case the context is that of polar (co-positive or co-negative definite, see Definition 2.5 below) meromorphic functions. In this setting, the determining measures are decomposed into principal and holomorphic parts, and a transition formula for $d\mu_{r}-d\mu_{l}$ was obtained under the assumption that the holomorphic components extend across the separatrix. A different point of view was considered by Harper, who established a connection between the existence of two Laplace representations for a holomorphic function $F$ on a vertical strip and solutions of the Cauchy problem for the heat equation on the infinite rod. By imposing suitable bounds on the $L^{2}$-norm of $F$ along vertical lines, Harper applied uniqueness results for the heat equation to obtain sufficient conditions under which $\mu_{r}=\mu_{l}$.
@@ 2 Notation and definitions
The classical bilateral Laplace transform
$\mathcal{L}(f)(s)=\int_{-\infty}^{+\infty}e^{-st}f(t)\,dt$
of a locally integrable real-valued function $f(t)$ may be seen as a special case of (1) by setting $\mu(t)=\int_{0}^{t}f(u)\,du.$ Note that this definition implies that $\mu(t)$ is normalized; we sometimes refer to $f(t)$ as normalized determining function and $(f_{l},f_{r})$ as Laplace pair, respectively. The usage will always be clear from the context.
@@
Note that co-positive definite functions correspond to classical positive definite functions on vertical lines within the strip; along each such line Widder’s characterization (2) reduces to Bochner’s theorem (see [26] for details).
@@
Let $F(s)=\int_{-\infty}^{+\infty}e^{-st}\,d\mu(t)$ have a vertical strip of convergence $\Omega_{a,b},{\rm with}-\infty<a<b<+\infty$. For every $\epsilon>0$ and $0<\delta<\frac{|b-a|}{2},$ there exists a positive constant $Y$ such that, for all $a+\delta\leq x\leq b-\delta$, ${\left|\frac{F(x+iy)}{y}\right|<\epsilon}$ whenever $|y|>Y.$
@@
We will need the version of the classical Phragmén-Lindelöf theorem for vertical half-strips.
@@
Let $D$ be a vertical half-strip defined by the points of $\Omega_{a,b}$ lying either above or below the line $\Im(s)=d.$ Let $F(s)$ be holomorphic in $D$ and continuous up to its boundary. Assume that $|F(s)|\leq M$ on boundary lines. If there exist positive constants $A,$ $B$ and $0<K<\frac{\pi}{|b-a|}$ such that for all $s$ inside $D$ we have
$|F(x+iy)|\leq Ae^{Be^{K|y|}},$
@@ 3 The transition formula
From this point onwards we suppose the function $F(s)$ is represented by a Laplace pair $(\mu_{l},\mu_{r})$ on a vertical strip $\Omega_{a,b}$ with $a<0<b$, where the imaginary axis $\Re(s)=0$ is the separatrix. This choice is notationally convenient and clearly implies no loss of generality, since a Laplace pair with separatrix $\Re(s)=c$ reduces to this setting after a translation of the determining measures by the factor $e^{-ct}$.
@@
Let $F(s)$ be a holomorphic Laplace pair $(\mu_{l},\mu_{r})$ satisfying the Phragmén-Lindelöf condition (4) and with no singularities on the separatrix. Let $0<\delta<\min\{|a|,|b|,1\}.$ Then there exist positive constants $R_{0}$ and $M$ such that
@@
By Lemma 2.8 it follows that $F(s)/s$ is bounded by some constant $M_{\delta}>0$ on the vertical lines $|\Re(s)|=\delta.$ Since $F(s)/s$ satisfies the Phragmén–Lindelöf condition, we may choose $R_{0}>0$ such that the bound (4) holds for $F(s)/s$ whenever $|y|\geq R_{0}.$
@@
We consider the contour integral ${\frac{1}{2\pi i}\int_{\Gamma_{R,\delta}}e^{st}\frac{F(s)}{s}\,ds.}$ To prove the result it suffices to show that, for a suitable choice of $\delta,$ the contributions of the upper and lower horizontal edges of $\Gamma_{R,\delta}$ become arbitrarily small for all sufficiently large $R.$ Indeed, under this assumption, the inversion formula (3) together with the residue theorem yields
$\lim_{R\to+\infty}\frac{1}{2\pi i}\int_{\Gamma_{R,\delta}}e^{st}\frac{F(s)}{s}\,ds=\mu_{r}(t)-\mu_{l}(t)+\mu_{l}(+\infty)-\mu_{r}(-\infty)=F(0).$ (6)
@@
As stated in the Introduction, the inversion argument used by Van der Pol [31] rested on the assumption that $F(s)/s$ vanishes uniformly inside the whole strip $\Omega_{a,b}$ as $|\Im(s)|\to+\infty.$ The Phragmén-Lindelöf condition (4) is the precise analytic mechanism that guarantees that this occurs. As will be shown in Section 6, if condition (4) is relaxed the growth of $F(s)$ on the separatrix may prevent the uniform vanishing of the horizontal edges of the contour, producing examples where $\mu_{r}\neq\mu_{l}.$
The following result generalizes Van der Pol’s transition formula to the case of a finite number of poles of arbitrary order.
@@
Let $F(s)$ be a meromorphic Laplace pair $(\mu_{l},\mu_{r})$ having finitely many poles on the imaginary axis $s=ip_{n}$, $1\leq n\leq N$, and satisfying the Phragmén–Lindelöf condition (4). Then for each real $t$
@@
Proceeding as in the proof of Theorem 3.2, the estimates (5) and (7) remain valid since none of the half-strips contain any poles by Remark 2.11. Therefore the same contour argument applies with the only modification that the poles $s=ip_{n}$ now lie inside the contour. By the residue theorem,
@@
Subtracting this identity from (9) yields (8), as claimed. ∎
@@
Let $F(s)$ be a meromorphic Laplace pair $(f_{l},f_{r})$ satisfying the hypotheses of Corollary 3.3. For each pole $s=ip_{n}$, let $r_{n}$ denote its order and let ${P_{n}(s)=\sum_{k=1}^{r_{n}}\frac{a_{n,k}}{(s-ip_{n})^{k}}}$ be the principal part of its Laurent expansion around $s=ip_{n}$. Then
$f_{r}(t)-f_{l}(t)=\sum_{n=1}^{N}e^{ip_{n}t}\sum_{k=1}^{r_{n}}a_{n,k}\frac{t^{k-1}}{(k-1)!}.$ (10)
@@
$e^{st}\frac{F(s)}{s}=e^{itp_{n}}\left[\sum_{\begin{subarray}{c}l,m\geq 0\\ k\geq-r_{n}\end{subarray}}\frac{(-1)^{m}}{(ip_{n})^{m+1}}\frac{t^{l}}{l!}b_{n,k}(s-ip_{n})^{m+l+k}\right],$
@@
Finally, summing over all poles yields (10), completing the proof ∎
Remark 3.5 (The Mellin Transform).
It is possible to rephrase our results in terms of the Mellin transform. Performing the change of variables $u=e^{-t}$ in the Laplace transform $\mathcal{L}(\mu)$, we obtain the Mellin transform
$\mathcal{M}(\beta)(s)=\int_{0}^{+\infty}t^{s}\,d\beta(t),\quad\beta(t)=-\mu(\log(1/t)).$
Similarly, for $f\in L^{1}_{\text{loc}}(\mathbb{R}),$ we obtain ${\mathcal{M}(g)(s)=\int_{0}^{+\infty}t^{s-1}g(t)\,dt},$ with $g(t)=f(\log(1/t)).$ Since the strip of convergence $\Omega_{a,b}$ remains unchanged, we define a Mellin pair $(\beta_{l},\beta_{r})$ or $(g_{l},g_{r})$ in a completely analogous manner to Definition 2.2. Indeed, the Mellin pair $(\beta_{l},\beta_{r})$ corresponds to the Laplace pair $(\mu_{l},\mu_{r})$ where $\mu_{r,l}(t)=-\beta_{r,l}(e^{-t})$ while the pair $(g_{l},g_{r})$ corresponds to $(f_{l},f_{r})$ with $f_{r,l}(t)=g_{r,l}(e^{-t}).$ This allows us to obtain the transition formulas corresponding to (8) and (10) in terms of Mellin pairs:
$\beta_{r}(t)-\beta_{l}(t)=\mu_{l}(\log(1/t))-\mu_{r}(\log(1/t))=\sum_{n=1}^{N}\operatorname*{Res}\bigg((1-t^{-s})\frac{F(s)}{s},\,s=ip_{n}\bigg);$
$g_{r}(t)-g_{l}(t)=\sum_{n=1}^{N}t^{-ip_{n}}\sum_{k=1}^{r_{n}}a_{n,k}\frac{(\log(1/t))^{k-1}}{(k-1)!}.$
@@ 4 The case of infinitely many poles
To establish the Laplace representation for $\Re(s)>0$, it suffices to show that for every $\epsilon>0$ there exists $C_{\epsilon}>0$ such that, for all $t\geq 0$, $|f(t)|\leq C_{\epsilon}e^{\epsilon t}.$
For $t\geq 0$, we have $\frac{t^{k-1}}{(k-1)!}\leq\frac{e^{\epsilon t}}{\epsilon^{k-1}},$ and therefore
@@
where $C_{\epsilon}<\infty$ by (14). This bound justifies the interchange of summation and integration, showing that $f(t)$ is locally integrable. For $\Re(s)>0$,
@@
Let the poles of $F(s)$ be such that its principal part $P(s)$ satisfies the hypotheses of Lemma 4.1. If the holomorphic part of $F$, defined by H(s)=F(s)P(s)H(s)=F(s)-P(s), satisfies the Phragmén–Lindelöf condition (4), then
@@
Since H(s)=F(s)P(s)H(s)=F(s)-P(s) is holomorphic and satisfies the Phragmén–Lindelöf condition, it follows from Theorem 3.2 that $\mu_{r,H}(t)=\mu_{l,H}(t)$, and therefore
@@
where $f(t)$ is given by the series (13). Estimate (15) now ensures that this series converges uniformly on compact sets, so that integration may be carried out termwise. Using the residue computations (11) and (12) in Lemma 3.4 for the finite pole case, we obtain the transition formula (16), as desired. ∎
@@ 5 Examples and applications
$\mu_{l}(t)=\begin{cases}\frac{1}{2}-\{e^{t}\},\,t>0,\,e^{t}\notin\mathbb{Z}\\ n+1-e^{t},\,t=\log(n+1),\,n\geq 0\\ \frac{1}{2}-e^{t},\,t<0\end{cases}$
A straightforward calculation shows that the maximal strip of convergence of $\mu_{l}$ is $\Omega_{0,1}$, so we denote it by $\mu_{0,1}$. Hence, on this strip, the Laplace representation
@@
We remark that in [5] (see section 4.5), the authors proved that $\zeta$ must lose its polarity upon crossing the separatrix $\Re(s)=1.$ Our explicit representation on $\Omega_{0,1}$ provides another proof of this fact: since $\mu_{0,1}$ is not monotonic, the corresponding measure is neither positive nor negative, implying that $F$ in neither co-positive definite nor co-negative definite – or, equivalently, that is possesses no polarity.
Since there are no poles on the imaginary axis, estimate (18) allows us to prove that the $\zeta$ function does not have a Laplace representation on any strip $\Omega_{a,0}$ adjacent to the left of the imaginary axis. Indeed, if that were the case, direct application of Theorem 3.2 shows that its determining measure would necessarily be equal to $\mu_{0,1}$. But this is impossible since $\Omega_{0,1}$ is the maximal strip of convergence of $\mu_{0,1}$.
@@
Observe that (20) and (22) recover the classical Laplace representations of $\frac{\zeta(s)}{s}$ on successive vertical strips of holomorphy as described in Titchmarsh [29].
@@
Let $f(t)$ be a bounded locally integrable periodic function of period $T$. Its Laplace transform
@@
The corresponding determining function is the $\pi$-periodic function $f(t)=\operatorname{sgn}(\sin t),$ whose Fourier series is given by ${\frac{2}{i\pi}\sum_{n\in\mathbb{Z}}\frac{1}{2n+1}e^{i(2n+1)t}.}$ While this series does not converge absolutely, it is not difficult to show that its Cèsaro means converge to $f$ in $L^{1}.$ Moreover, this is always true for bounded periodic functions. Therefore, when $f(t)$ is a bounded locally integrable periodic determining function, the transition formula (23) can always be interpreted as converging in $L^{1}$ via Cèsaro means.
@@
Example 5.4 (A quotient of $\Gamma$ functions).
@@
$_{1}F_{1}(a:b;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}}{(b)_{n}}\frac{z^{n}}{n!},\qquad{\rm where}\ (a)_{n}=\frac{\Gamma(a+n)}{\Gamma(a)},$
@@
Stirling’s formula for the $\Gamma$ function (see [25], 2.1.8) shows that